Verify that y=e^2xsinx is a solution to the differential equation (d^2/dx^2)-(4dy/dx)+5y=0


(2) Differentiate ln(1+sin^2x)

(3) differentiate x^x
(4).Evaluate the integral [ (2x^3-4x-8)/(x^4-x^3+4x^2-4x)]dx
1
Expert's answer
2020-06-29T19:24:22-0400

Given "y=e^{2x}sinx \implies y' = e^{2x}cosx+2e^{2x}sinx"

"\implies y''-4y'+5y = e^{2x}(4cosx+3sinx-4cosx-8sinx+5sinx)
= e^{2x} (0) = 0"

Hence, "y =e^{2x} sinx" is solution of "y''-4y'+5y=0" .


2) Let "y = ln(1+sin^2x)"

So, by differentiation with respect to x, we get

"\frac{dy}{dx} = \frac{1}{1+sin^2x} (2sinxcosx)" .


3) Let "y = x^x" "\implies log(y)= xlog(x)"

So, by differentiation with respect to x, we get

"\frac{1}{y}\frac{dy}{dx} = 1+log(x)
\implies \frac{dy}{dx} = x^x(1+logx)" .


4) We partial fraction,

"\frac{\left(2x^3-4x-8\right)}{\left(x^4-x^3+4x^2-4x\right)} = \frac{2}{x}-\frac{2}{x-1}+\frac{2x+4}{x^2+4}" = "\frac{2}{x}-\frac{2}{x-1}+\frac{2x}{x^2+4} + \frac{4}{x^2+4}"

Hence, "\intop\frac{\left(2x^3-4x-8\right)}{\left(x^4-x^3+4x^2-4x\right)} dx= \intop( \frac{2}{x}-\frac{2}{x-1}+\frac{2x}{x^2+4} + \frac{4}{x^2+4})dx"

= "2ln(x)-2ln(x-1)+ln(x^2+4)+ 2tan^{-1}(\frac{x}{2}) +c="

"= ln(\frac{x^2(x^2+4)}{(x-1)^2}) +2tan^{-1}(\frac{x}{2})+c" .